OpenStudy (anonymous):

Evaluate (5 + 2)^2 − 40 + 3

jimthompson5910 (jim_thompson5910):

Any ideas on the first step?

jimthompson5910 (jim_thompson5910):

Using PEMDAS, the first thing do to is to evaluate everything in parenthesis

jimthompson5910 (jim_thompson5910):

So 5+2 = 7 which means (5 + 2)^2 − 40 + 3 becomes (7)^2 − 40 + 3

jimthompson5910 (jim_thompson5910):

What's next

OpenStudy (anonymous):

i just stared goin back to school after 5 years so i dont know much

jimthompson5910 (jim_thompson5910):

7^2 = 7*7 = what?

OpenStudy (anonymous):

49

jimthompson5910 (jim_thompson5910):

so (7)^2 − 40 + 3 becomes 49 − 40 + 3

jimthompson5910 (jim_thompson5910):

do you see where to go from here?

OpenStudy (anonymous):

12

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

the final answer is 12

OpenStudy (anonymous):

thank u

jimthompson5910 (jim_thompson5910):

np